OpenStudy (anonymous):

The gravitational force between two spheres 5 meters apart from each other is 48 Newtons. How many Newtons of force will result when the distance between the two spheres changed from its original distance of 5 meters apart to 25 meters apart?

OpenStudy (whpalmer4):

This can be done a bit more simply by realizing from the equation given that the force varies inversely with the square of the distance between them, so \[F = \frac{k}{r^2}\]We know that when \(r = 5\), \(F = 48\), so \[48 = \frac{k}{(5)^2}\]or\[k = 48*25\]Now we want \(F\) when \(r = 25\), so \[F = \frac{48*25}{(25)^2} = \frac{48}{25} = 1.92\] Or another route, \[F = \frac{k}{r^2}\]We are increasing \(r\) by a factor of \(5\), so replace \(r\) by \(5r\): \[F_{new} = \frac{k}{(5r)^2} = \frac{k}{25r^2}\]and it should be evident upon comparison that that is \(1/25\) of the original force, or \(0.04 * 48 \text{ N} = 1.92 \text{ N}\)